### v1

# Certain Inequalities for the generalized polar derivative of a polynomial

Abstract.
Recently Rather et al. [9] considered the generalized derivative and the generalized polar derivative and studied the relative position of zeros of generalized derivative and generalized polar derivative with respect to the zeros of polynomial.
In this paper, we establish some inequalities that estimate the maximum modulus of generalized derivative and the generalized polar derivative of the polynomial $P(z)$, which is also the extension of recently known results.
Keywords: Polynomials, Generalized polar derivative, Maximum modulus of $P(z)$.
2010 Mathematics Subject Classification: 26D10, 41A17, 30A10, 26D05
1. Introduction
Let $P(z)$ be a polynomial of degree $n$. A well known result, known as Bernstein’s inequality [14], provides an upper bound for the maximum modulus of $P^{\prime}(z)$ on the unit circle $|z|=1$ in terms of the maximum modulus of $P(z)$, which is given by
$\displaystyle\max_{|z|=1}|P^{\prime}(z)|\leq n\displaystyle\max_{|z|=1}|P(z)|.$ (1)
Equality holds if and only if all the zeros of $P(z)$ are at the origin. For polynomials whose zeros lie entirely within the unit disk $|z|\leq 1$, Turán [15] established a lower bound for the maximum modulus of $P^{\prime}(z)$ on $|z|=1$. Specifically, if $P(z)$ is a polynomial of degree $n$ with all zeros in $|z|\leq 1$, then
$\displaystyle\max_{|z|=1}|P^{\prime}(z)|\geq\frac{n}{2}\displaystyle\max_{|z|=1}|P(z)|.$ (2)
This inequality is sharp, and equality holds for $P(z)=\alpha z^{n}+\beta$, where $|\alpha|=|\beta|\neq 0$.
Now for the estimation of the lower bounds of $Re\Big(\frac{zP^{\prime}(z)}{P(z)}\Big)$ on $|z|=1$, Dubinin [5] proved the following.
Theorem A.
If $P(z)\in P_{n}$ has all its zeros in $|z|\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
$Re\Big(z\frac{P^{\prime}(z)}{P(z)}\Big)\geq\frac{n}{2}\Big\{1+\frac{1}{n}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big\}.$ (3)
As an application of this, Dubinin [5] obtained an interesting refinement of (2), by proving that if all the zeros of $P(z)$ lie in $|z|\leq 1$, then
$\displaystyle\max_{|z|=1}|P^{\prime}(z)|\geq\frac{1}{2}\Big(n+\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\displaystyle\max_{|z|=1}|P(z)|.$ (4)
Malik [7] extended $(2)$ and proved that if $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$, then
$\underset{|z|=1}{max}|P^{\prime}(z)|\geq\frac{n}{1+k}\underset{|z|=1}{max}|P(z)|.$ (5)
Equality in (5) holds for the polynomial $P(z)=(z+k)^{n}$.
Rather et al. [12] generalized Theorem A and proved the following.
Theorem B.
If $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
$Re\Big(z\frac{P^{\prime}(z)}{P(z)}\Big)\geq\frac{n}{1+k}\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\},$ (6)
which in turn yields the following refinement of the inequality $(5)$ as well as a generalization of the inequality $(4)$
Theorem C.
If $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for $|z|=1$,
$|P^{\prime}(z)|\geq\frac{n}{1+k}\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\}|P(z)|.$ (7)
Let $D_{\alpha}[P](z)$ denote the polar differentiation [8] of a polynomial $P(z)$ of degree $n$ with respect to a complex number $\alpha$, then
$D_{\alpha}[P](z)=nP(z)+(\alpha-z)P^{\prime}(z).$
Note that the polynomial $D_{\alpha}[P](z)$ is of degree at most n1n-1 and it generalizes the ordinary derivative $P^{\prime}(z)$ of $P(z)$ in the sense that
$\underset{\alpha\rightarrow\infty}{\lim}\frac{D_{\alpha}[P](z)}{\alpha}=P^{\prime}(z)$
uniformly with respect $z$ for $\left|z\right|\leq R,R>0.$
The Bernstein-type inequalities for the class of polynomials with “ordinary derivative” replaced by “polar derivative” have attracted number of mathematicians. In this direction, Aziz [2] was the first to establish inequalities concerning the polar derivative of a polynomial in terms of the modulus of the polynomial on the unit disk.
As an extension of the inequality $(1)$ to the polar derivative, Aziz [2] proved the following result.
If $P(z)$ is a polynomial of degree n, then for every $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$ we have
$|D_{\alpha}[P](z)|\leq n|\alpha z^{n-1}|~\underset{|z|=1}{max}|P(z)|\quad for\quad|z|\geq 1.$
The result is best possible and equality in the above inequality holds for $P(z)=cz^{n},\quad c\neq 0.$
Concerning the class of polynomials having all zeros in $|z|\leq k$, Aziz and Rather [3] obtained several sharp results concerning the maximum modulus of $D_{\alpha}[P](z)$ on $|z|=1.$ Among other things, they established the following extension of the inequality (5) to the polar derivative of a polynomial.
Theorem D.
If all the zeros of P(z) lie in $\left|z\right|\leq k$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k\leq 1,$
$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq\frac{n}{1+k}(|\alpha|-k)\underset{|z|=1}{max}|P(z)|.$ (8)
Rather et al. [12] refined this inequality by proving the following result
Theorem E.
If all the zeros of $P(z)$ lie $|z|\leq k,k\leq 1$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
$\displaystyle\max_{|z|=1}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k}\Big)\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\}\displaystyle\max_{|z|=1}|P(z)|.$ (9)
In literature there exist several generalizations and refinements of these inequalities. For $P(z)=c(z-z_{1})(z-z_{2})\cdots(z-z_{n})$ and $\mathbb{R}_{+}^{n}$ being the set of all n-tuples $\gamma:=(\gamma_{1},\gamma_{2},\ldots,\gamma_{n})$ of non-negative real numbers (not all zeros) with $\sum\limits_{j=1}^{n}\gamma_{j}=\wedge$ , Sz-Nagy [13] introduced generalized derivative of polynomial $P(z)$ defined by
$P^{\gamma}(z):=P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}=\sum\limits_{j=1}^{n}\gamma_{j}P_{k}(z).$
Notice that for $\gamma=(1,1,1,\ldots,1),P^{\gamma}(z)=P^{\prime}(z).$ In view of this, we call it generalized derivative of polynomial $P(z)$.
Next we define generalized polar derivative of $P(z)$ as
$D_{\alpha}^{\gamma}[P](z):=\wedge P(z)+(\alpha-z)P^{\gamma}(z),$
where $\wedge=\sum\limits_{j=1}^{n}\gamma_{j}.$
Noting that for $\gamma=(1,1,1,\ldots,1),\quad D_{\alpha}^{\gamma}[P](z)=D_{\alpha}P(z)$, we call it generalized polar derivative of $P(z)$.
Rather et al. [11] extended Theorem D to the generalized polar derivative and proved the following result.
Theorem F.
If all the zeros of P(z) lie in $\left|z\right|\leq k$, $k\leq 1$ then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\frac{\wedge}{1+k}(|\alpha|-k)\underset{|z|=1}{max}|P(z)|.$
Recently Rather et al. [9] proved the following results
Theorem G.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then
$|P^{\gamma}(z)|\geq\frac{k}{1+k}\Big[\frac{\wedge}{k}+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]|P(z)|,$ (10)
where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$.
Theorem H.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k}\Big)\Big[\wedge+k\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$ and the result is best possible for polynomial $P(z)=(z+k)^{n}$
For $k=1$, Theorem H reduces to the following result.
Theorem I.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq 1$, then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-1}{2}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$ and the result is best possible for polynomial $P(z)=(z+1)^{n}$.
2. Main Results
[…]
Theorem 1.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|P^{\gamma}(z)|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$ (11)
where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$. The result is best possible for polynomial $P(z)=(z+k)^{n}$
If we take $\gamma=(1,1,\cdots,1)$ in Theorem 11, we get the following result.
Corollary 1.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|P^{\prime}(z)|\geq\frac{n}{1+k^{n}}\Big[1+\frac{1}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$ (12)
Theorem 2.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$.
If we take $\gamma=(1,1,\cdots,1)$ in Theorem 22, we get the following result.
Corollary 2.
If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[1+\frac{1}{n}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
For the proof of the theorems we need following lemmas, the first is due to Rather et al. [12].
Lemma 1.
If $0\leq x_{j}\leq 1$,    $j=1,2,\cdots,n,$ then
$\displaystyle\sum_{j=1}^{n}\frac{1-x_{j}}{1+x_{j}}\geq\frac{1-\prod_{j=1}^{n}x_{j}}{1+\prod_{j=1}^{n}x_{j}}\qquad\forall n\in N.$
From a simple consequence of Maximum Modulus theorem, we can get the following result.
Lemma 2.
If $P(z)$ is a polynomial of degree n, then for $R\geq 1$
$\underset{|z|=R}{max}|P(z)|\leq R^{n}\underset{|z|=1}{max}|P(z)|.$
Lemma 3.
[3]. If $P(z)\in\mathcal{P}_{n},$ having all its zeros in the disk $|z|\leq k$ where $k\geq 1,$ then
$\underset{|z|=k}{max}|P(z)|\geq\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$ (13)
3. Proof of Theorems
Proof of Theorem 1.
Since $P(z)$ is a polynomial of degree $n$ has all its zeros in $|z|\leq k$, $k\geq 1$.
Let $G(z)=P(kz)$, then all the zeros of $G(z)$ lie in $|z|\leq 1$.
Now
$G^{\gamma}(z)=G(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-w_{j}},$
where $w_{j}=\frac{z_{j}}{k}$ and $|w_{j}|\leq 1$.
For $|z|=1,$
$|zG^{\gamma}(z)|=$ $|G(z)|\Big|\displaystyle\sum_{j=1}^{n}\frac{z\gamma_{j}}{z-w_{j}}\Big|$
$\geq$ $|G(z)|\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{1+|w_{j}|}$
$=$ $\Big\{\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{2}\Big(1+\frac{1-|w_{j}|}{1+|w_{j}|}\Big)\Big\}|G(z)|$
$=$ $\Big(\frac{\wedge}{2}+\frac{1}{2}\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}(1-|w_{j}|)}{1+|w_{j}|}\Big)|G(z)|$
$\geq$ $\Big(\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\displaystyle\sum_{j=1}^{n}\frac{1-|w_{j}|}{1+|w_{j}|}\Big)|G(z)|$
$=$ $\Big(\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\displaystyle\sum_{j=1}^{n}\frac{1-\frac{|z_{j}|}{k}}{1+\frac{|z_{j}|}{k}}\Big)|G(z)|$
$\geq$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{1-\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}{1+\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}\Big)\Big]|G(z)|$
$=$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]|G(z)|.$
This mplies
$\displaystyle\max_{|z|=1}\Big|G(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-w_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|G(z)|$
$\text{and so}\displaystyle\max_{|z|=1}\Big|P(kz)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(kz)|$
Now from Lemma 33, we have
$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
$Thus\qquad k\displaystyle\max_{|z|=k}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
$so\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(z)|$
This proves the theorem.
Proof of Theorem 2.
Since the polynomial $P(z)$ has all its zeros in $|z|\leq k,$ where $k\geq 1.$
Hence the polynomial $G(z)=P(kz)$ has all its zeros in $|z|\leq 1$.
Applying Theorem I to the polynomial $G(z)$ and noting that $\big|{\frac{\alpha}{k}}\big|\geq 1,$ we get
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[G](z)\bigg|\geq$ $\frac{\frac{|\alpha|}{k}-1}{2}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|.$
Replacing $G(z)$ by $P(kz)$, we get
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(kz)|$
$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
With the help of Lemma 33, we have
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
$\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$
This gives
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$ (14)
Also we have
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $=\underset{|z|=1}{max}\bigg|\wedge P(kz)+\bigg(\frac{\alpha}{k}-z\bigg)P^{\gamma}(kz)\bigg|$
$=\underset{|z|=1}{max}\bigg|\wedge P(kz)+\bigg(\frac{\alpha-kz}{k}\bigg)P(kz)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\bigg|$
$=\underset{|z|=1}{max}\bigg|\wedge P(kz)+\bigg(\frac{\alpha-kz}{k}\bigg)kP(kz)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{kz-z_{j}}\bigg|$
$=\underset{|z|=1}{max}\bigg|\wedge P(kz)+(\alpha-kz)P(kz)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{kz-z_{j}}\bigg|$
$=\underset{|z|=1}{max}|F(kz)|$
$=\underset{|z|=k}{max}|F(z)|,$
where $F(z)=\wedge P(z)+(\alpha-z)P(z)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}$ is a polynomial of degree atmost n1n-1.
On using Lemma 22, this gives
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $=\underset{|z|=k}{max}|F(z)|$
$\leq k^{n-1}\underset{|z|=1}{max}|F(z)|$
$=k^{n-1}\underset{|z|=1}{max}\bigg|\wedge P(z)+(\alpha-z)P(z)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\bigg|$
$=k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|.$
Hence,
$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|\leq k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|.$ (15)
Combining the inequalities $(14)$ and $(15)$, we get
$k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$
This implies,
$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
This proves the result. ∎

### v2

@@
 This inequality is sharp, and equality holds for $P(z)=\alpha z^{n}+\beta$, where $|\alpha|=|\beta|\neq 0$.
-Now for the estimation of the lower bounds of $Re\Big(\frac{zP^{\prime}(z)}{P(z)}\Big)$ on $|z|=1$, Dubinin [5] proved the following.
+The estimation of the lower bounds of $Re\Big(\frac{zP^{\prime}(z)}{P(z)}\Big)$ on $|z|=1$, Dubinin [5] proved the following.
 Theorem A.
-If $P(z)\in P_{n}$ has all its zeros in $|z|\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ has all its zeros in $|z|\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
 $Re\Big(z\frac{P^{\prime}(z)}{P(z)}\Big)\geq\frac{n}{2}\Big\{1+\frac{1}{n}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big\}.$ (3)
@@
 Theorem B.
-If $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ has all its zeros in $|z|\leq k$, $k\leq 1$, then for all $z$ on $|z|=1$, for which $P(z)\neq 0$,
 $Re\Big(z\frac{P^{\prime}(z)}{P(z)}\Big)\geq\frac{n}{1+k}\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\},$ (6)
@@
 Theorem C.
-If $P(z)\in P_{n}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for $|z|=1$,
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ has all its zeros in $|z|\leq k$, $k\leq 1$ then for $|z|=1$,
 $|P^{\prime}(z)|\geq\frac{n}{1+k}\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\}|P(z)|.$ (7)
@@
 In this direction, Aziz [2] was the first to establish inequalities concerning the polar derivative of a polynomial in terms of the modulus of the polynomial on the unit disk.
-As an extension of the inequality $(1)$ to the polar derivative, Aziz [2] proved the following result.
+As an extension of the inequality $(1)$ to the polar derivative, Aziz [2] proved that
 if $P(z)$ is a polynomial of degree n, then for every $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$ we have
@@
 Theorem E.
-If all the zeros of $P(z)$ lie $|z|\leq k,k\leq 1$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
+If all the zeros of $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ lie $|z|\leq k,k\leq 1$, then for $\alpha\in\mathbb{C}$ with $\left|\alpha\right|\geq k,$
 $\displaystyle\max_{|z|=1}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k}\Big)\Big\{1+\frac{k}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big\}\displaystyle\max_{|z|=1}|P(z)|.$ (9)
@@
 Theorem G.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then
 $|P^{\gamma}(z)|\geq\frac{k}{1+k}\Big[\frac{\wedge}{k}+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]|P(z)|,$ (10)
@@
 Theorem H.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\leq 1,$ then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
 $\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k}\Big)\Big[\wedge+k\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
@@
 Theorem I.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq 1$, then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq 1$, then for $\alpha\in\mathbb{C}$ with $|\alpha|\geq 1,$
 $\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-1}{2}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
@@
 Theorem 1.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
 $\underset{|z|=1}{max}|P^{\gamma}(z)|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$ (11)
 where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$. The result is best possible for polynomial $P(z)=(z+k)^{n}$
-If we take $\gamma=(1,1,\cdots,1)$ in Theorem 11, we get the following result.
+If we take $\gamma=(1,1,\cdots,1)$ in Theorem $1$, we get the following result.
 Corollary 1.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
 $\underset{|z|=1}{max}|P^{\prime}(z)|\geq\frac{n}{1+k^{n}}\Big[1+\frac{1}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$ (12)
 Theorem 2.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
-$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
+$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|,$ (13)
 where $\gamma_{m}=\min\{\gamma_{1},\gamma_{2},\cdots,\gamma_{n}\}$.
-If we take $\gamma=(1,1,\cdots,1)$ in Theorem 22, we get the following result.
+Remark 1.
+If we divide inequality (13) by $|\alpha|$ and then letting $|\alpha|\rightarrow\infty$, we get the inequality
+$\underset{|z|=1}{max}|P^{\gamma}(z)|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$ (14)
+If we take $\gamma=(1,1,\cdots,1)$ in Theorem $2$, we get the following result.
 Corollary 2.
-If $P(z)$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
-$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[1+\frac{1}{n}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
-For the proof of the theorems we need following lemmas, the first is due to Rather et al. [12].
+If $P(z)=\displaystyle\sum_{j=0}^{n}a_{j}z^{j}$ is a polynomial of degree $n$ having all its zeros in $\left|z\right|\leq k$ where $k\geq 1,$ then
+$\underset{|z|=1}{max}|D_{\alpha}[P](z)|\geq n\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[1+\frac{1}{n}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
+For the proof of the theorems we need following lemmas. The first is due to Rather et al. [12].
 Lemma 1.
@@
 $\displaystyle\sum_{j=1}^{n}\frac{1-x_{j}}{1+x_{j}}\geq\frac{1-\prod_{j=1}^{n}x_{j}}{1+\prod_{j=1}^{n}x_{j}}\qquad\forall n\in N.$
-From a simple consequence of Maximum Modulus theorem, we can get the following result.
+By a simple consequence of Maximum Modulus theorem, we can get the following result.
 Lemma 2.
@@
 Lemma 3.
-[3]. If $P(z)\in\mathcal{P}_{n},$ having all its zeros in the disk $|z|\leq k$ where $k\geq 1,$ then
+[3]. If $P(z)$ is a polynomial having all its zeros in the disk $|z|\leq k$ where $k\geq 1,$ then
 $\underset{|z|=k}{max}|P(z)|\geq\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$ (15)
@@
 Proof of Theorem 1.
-Since $P(z)$ is a polynomial of degree $n$ has all its zeros in $|z|\leq k$, $k\geq 1$.
+Since $P(z)$ is a polynomial of degree $n$ having all its zeros in $|z|\leq k$, $k\geq 1$.
 Let $G(z)=P(kz)$, then all the zeros of $G(z)$ lie in $|z|\leq 1$.
@@
 $=$ $\Big(\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\displaystyle\sum_{j=1}^{n}\frac{1-\frac{|z_{j}|}{k}}{1+\frac{|z_{j}|}{k}}\Big)|G(z)|$
-$\geq$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{1-\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}{1+\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}\Big)\Big]|G(z)|$
+Using Lemma 1, we get,
+$|zG^{\gamma}(z)|\geq$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{1-\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}{1+\displaystyle\prod_{j=1}^{n}\frac{|z_{j}|}{k}}\Big)\Big]|G(z)|$
 $=$ $\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]|G(z)|.$
@@
 $\displaystyle\max_{|z|=1}\Big|G(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-w_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|G(z)|$
-$\text{and so}\displaystyle\max_{|z|=1}\Big|P(kz)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(kz)|$
-Now from Lemma 33, we have
-$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
-$Thus\qquad k\displaystyle\max_{|z|=k}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|$
-$so\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(z)|$
+Replacing $G(z)$ by $P(kz)$, we get,
+$\displaystyle\max_{|z|=1}\Big|P(kz)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-\frac{z_{j}}{k}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(kz)|$
+$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=k}|P(z)|$
+Now from Lemma $3$, we have
+$k\displaystyle\max_{|z|=k}\Big|P(z)\displaystyle\sum_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
+Thus we have,
+$k\displaystyle\max_{|z|=k}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
+Using Lemma 2, we get,
+$k^{n}\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\Big[\frac{\wedge}{2}+\frac{\gamma_{m}}{2}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\displaystyle\max_{|z|=1}|P(z)|.$
+That is,
+$\displaystyle\max_{|z|=1}\Big|P^{\gamma}(z)\Big|\geq\frac{\wedge}{1+k^{n}}\Big[1+\frac{\gamma_{m}}{\wedge}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\displaystyle\max_{|z|=1}|P(z)|$
 This proves the theorem.
 Proof of Theorem 2.
-Since the polynomial $P(z)$ has all its zeros in $|z|\leq k,$ where $k\geq 1.$
-Hence the polynomial $G(z)=P(kz)$ has all its zeros in $|z|\leq 1$.
+Since the polynomial $P(z)$ has all its zeros in $|z|\leq k$ where $k\geq 1.$
+Therefore the polynomial $G(z)=P(kz)$ has all its zeros in $|z|\leq 1$.
 Applying Theorem I to the polynomial $G(z)$ and noting that $\big|{\frac{\alpha}{k}}\big|\geq 1,$ we get
-$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[G](z)\bigg|\geq$ $\frac{\frac{|\alpha|}{k}-1}{2}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|$
-$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|.$
+$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[G](z)\bigg|\geq$ $\frac{\frac{|\alpha|}{k}-1}{2}\Big[\wedge+\gamma_{m}\Big(\frac{|k^{n}a_{n}|-|a_{0}|}{|k^{n}a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|$
+$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|G(z)|.$
 Replacing $G(z)$ by $P(kz)$, we get
-$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(kz)|$
-$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
-With the help of Lemma 33, we have
-$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
-$\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$
+$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(kz)|$
+$=$ $\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
+With the help of Lemma $3$, we have
+$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=k}{max}|P(z)|$
+$\geq\frac{|\alpha|-k}{2k}\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\frac{2k^{n}}{1+k^{n}}\underset{|z|=1}{max}|P(z)|.$
 This gives
-$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$ (14)
+$\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$ (16)
 Also we have
@@
 where $F(z)=\wedge P(z)+(\alpha-z)P(z)\sum\limits_{j=1}^{n}\frac{\gamma_{j}}{z-z_{j}}$ is a polynomial of degree atmost n1n-1.
-On using Lemma 22, this gives
+On using Lemma $2$, this gives
 $\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|$ $=\underset{|z|=k}{max}|F(z)|$
@@
 $\underset{|z|=1}{max}\bigg|D_{\frac{\alpha}{k}}^{\gamma}[P](kz)\bigg|\leq k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|.$ (17)
-Combining the inequalities $(14)$ and $(15)$, we get
-$k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$
+Combining inequalities $(14)$ and $(15)$, we get
+$k^{n-1}\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]k^{n-1}\underset{|z|=1}{max}|P(z)|.$
 This implies,
-$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{|a_{n}|-|a_{0}|}{|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
+$\underset{|z|=1}{max}|D_{\alpha}^{\gamma}[P](z)|\geq\Big(\frac{|\alpha|-k}{1+k^{n}}\Big)\Big[\wedge+\gamma_{m}\Big(\frac{k^{n}|a_{n}|-|a_{0}|}{k^{n}|a_{n}|+|a_{0}|}\Big)\Big]\underset{|z|=1}{max}|P(z)|.$
 This proves the result. ∎